OpenStudy (anonymous):

When you have a problem like 3(x+1/2)-1/3x+11 what are the steps? I came up with 3x+3_1/2-1/3x+11 and then simplified to 2_2/3x+14_1/2 and simplified that to 8/3x+29/2. The question asks me to simplify the expression if I can. If it cannot be simplified then express so.

OpenStudy (anonymous):

is it \[3(x+\frac{1}{2})-\frac{1}{3}x+11\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

first step is to multiply using distributive property

OpenStudy (anonymous):

\[3x+\frac{3}{2}-\frac{1}{3}x+11\] then combine like terms

OpenStudy (anonymous):

I did that. But it keeps telling me it is wrong when I simplify all the way down.

OpenStudy (anonymous):

but \(3\times \frac{1}{2}=\frac{3}{2}=1\tfrac{1}{2}\) not \(3\tfrac{1}{2}\)

OpenStudy (anonymous):

3_1/2x

OpenStudy (anonymous):

lets forget about mixed numbers

OpenStudy (anonymous):

\[3x+3\frac{ 1 }{ 2 }-\frac{ 1 }{ 3 }x+11\]

OpenStudy (anonymous):

that is a mistake

OpenStudy (anonymous):

Combined terms would equal \[2\frac{ 2 }{ 3 }+14\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

forgot the x on the first part.

OpenStudy (anonymous):

lets try again parentheses mean multiplication, not addition so \[3(x+\frac{1}{2}=3\times x+3\times \frac{1}{2}=3x+\frac{3}{2}\]

OpenStudy (anonymous):

you have \(3\tfrac{1}{2}\)and it is a mistake, it should be \(\frac{3}{2}\)

OpenStudy (anonymous):

I see what I did wrong I think... it should come out to be \[\frac{ 8 }{ 3 }x+\frac{ 25 }{ 2 }\]

OpenStudy (anonymous):

I still don't see how you are coming up with 3/2

OpenStudy (anonymous):

3x-1/3x = 2_2/3x then 1_1/2 + 11 = 12_1/2

OpenStudy (anonymous):

I found the correct answer now. Thanks for the reminder!